Math 110:Basic Calculus1 Fall 1998

SYLLABUS

Ron Buckmire Ramin Naimi

Where and When: Fowler 316, 8^{30} am−9^{25} am and
Fowler 316, 9^{30} am−10^{25} am

Texts: Calculus in Context by Callahan and Hoffman, Calculus by Hughes-Hallet, Gleason et al. (2nd Edition)

Instructors:

Ron Buckmire: Of�ce Hours: M 10^{30} am−11^{30} am, Fowler 320
W 1^{30} pm−3^{30} pm

Phone: x2536

email: buckmire@oxy.edu

Ramin Naimi: Of�ce Hours: MWF 1^{30} pm−2^{30} pm, Fowler 403

Phone: x2550

email: rnaimi@oxy.edu

Important Course Information:

Course Description: Models in the natural and social sciences often take the form of a system of differential equations. We approach the study of calculus by examining how these models are constructed and used for analysis. The mathematical theme running throughout the course is that of local linearity and error estimation. The chief goals of this course are to understand and master the techniques associated with differential calculus, increase analytical skills, improve communication and writing skills in mathematics and learn to love math or at least to appear to love math.

Homework: Homework will be assigned daily, collected weekly and graded. You are, of course, strongly encouraged to do at least all of the assigned problems. In order to learn mathematics, one must practice mathematics. You are encouraged to work in groups, but you must turn in your own homework.

Quizzes: During the semester, there will be approximately 9 quizzes. They will be usually takehome quizzes and will resemble homework problems which connect or advance important concepts and skills. Work on quizzes are subject to the same rules as exams.

Labs and Lab Write-ups: Labs are not optional. In the labs, you will have the chance to collaborate with your fellow students. It is your opportunity to struggle with the dif�cult skill of successfully communicating your ideas and actively listening to the ideas of others as you work together to produce well-written lab reports and essays.

In-ClassExams: There will be three in-class exams. Exams will be given on Thursday September 24 and Thursday October 29 (in your lab section) and Monday November 23.

Final Exam: TBA. As soon as we know when our �nal will be held, we will pass the information on to you. Do not set a date for your return home for the holidays until you KNOW when you will be �nished with your �nal exams. The rules of Occidental College will not recognize travel plans as an excuse to move a �nal exam time.

Gateway Exams: There will be four (4) Gateway Exams that will be given this semester. Failure to pass a Gateway will result in lowering your �nal grade one third of a grade. See the handout for more information on the Gateway exams.

Independent and Group Study: Students are strongly encouraged to work in groups, although work turned in for evaluation must be your own.

Online Component: We have class mailing lists, to which all registered students in Math110 are subscribed. The addresses are Math110sec1-L@oxy.edu and Math110sec2-L@oxy.edu. You should use this mailing list to communicate with your colleagues in the class in a professional manner. In addition, we have a class website, available at http://abacus.oxy.edu/ron/math/110/98/ . You can use the class website to review homework solutions, check on the course schedule and give feedback to professors on the course.

Calculator Use: You will need at least a scienti�c calculator with graphing capabilities for this course. If you do not at present own a calculator, we recommend you buy the TI-83 graphing calculator. If you already own a different graphing calculator, you do not need to go out and buy a TI-83, but you may have some work to do to �gure out how our TI-83 speci�c instructions can be translated to work with your calculator. We will be usign the TI-83 in class, in lab and on exams. The use of calculators on exams is encouraged, BUT inappropriate use will not be tolerated. For instance, using the programming capabilties to record notes is dishonest work. If a question on an exam expressly forbids the use of the graphing capabilities of your calculator, it means just that. If you have any doubt about using any features on your calculator on exams, ask one of the instructors. Do not trust your classmates to know what is allowed and what is not allowed. If you are caught using your calculator in an unacceptable manner, your actions will be reported to the proper authorities.

Course Policies: You are expected to know and follow the policies below.

Honest Academic Work: It is expected that each student in this class will conduct her- or himself within the guidelines of the Student Handbook. All academic work should be done with the complete honesty and integrity that this college demands. Classroom Conduct: Our primary goal in this classroom is to teach/learn/discuss/debate/enjoy/do calculus. This is best accomplished when we feel free to question and doubt, free to argue and exchange creative ideas. If one feels threatened or unwelcome, this becomes impossible. Therefore, the classroom should be a safe space. All are welcomed and encouraged to actively participate in the learning of calculus, regardless of gender, race, nationality, native language, sexuality, political ideology, and especially personal mathematical history. Any student who feels she or he is experiencing a hostile environment should speak to one of us.

Make-Up Work: No late work will be accepted this semester. No make-up quizzes or exams will be given. If you must miss a quiz or exam, let us know as soon as possible beforehand. If we feel your excuse is reasonable, we will work something out. If you do not alert us in advance of the possibility of missed work, there is no chance for an accomodation to be made.

Tardiness: There is no excuse for being late for lectures or labs. Entering late disrupts the �ow of class and sends the message that you do not respect your fellow students or your professors. If you arrive late, enter quietly and deal with missed handouts after class. If you will be late on a regular basis, please come and share the reasons with us before we approach you.

Grading: The table below explains how your �nal average in this course will be determined.

Score  percent
Homework Average  × 10%  =
Quiz Average  × 10%  =
Exam � 1  × 15%  =
Exam � 2  × 15%  =
Exam � 3  × 15%  =
Lab Average  × 15%  =
Final Exam  × 20%  =

Total:

The grade scale below assumes that you have passed all of the gateway exams:

average {below \atop 60}  60�68  68�70  70�72  72�78  78�80
letter grade F  D  D+  C-  C  C+
average 80�82  82�88  88�90  90-92  93 or above  100
letter grade B-  B  B+  A-  A  A(+)

Unfortunately, there is no of�cial A+. You are expected to keep track of your own quiz and lab averages. You should also keep all of your quizzes and labs. This way, discrepancies can be worked out easily.